### v1

# Optimal Trudinger-Moser inequalities on complete noncompact Riemannian manifolds: Revisit of the argument from the local inequalities to global ones

Abstract.
The main purpose of this short note, on the one hand, to is clarify some part of the proof of Theorem 1.3 in [8] in a simple way, and on the other hand, to give an alternative argument from local inequalities to global ones.
1. Introduction
The classical Pólya-Szegö symmetrization inequality plays a fundamental role in establishing geometric and functional inequalities in the Euclidean space. However, this symmetrization principle in general fails in many other non-Euclidean settings such as the Heisenberg group or Riemannian manifolds as well as on higher order Sobolev spaces even in the Euclidean space.
To circumvent this obstacle, Lam and Lu [5, 6] developed a symmetrization-free method to prove critical global Trudinger-Moser inequalities on the entire Heisenberg group, and Adams inequalities on the higher order Sobolev spaces on the entire Euclidean spaces. Lam, Lu and Tang [7] establish the subcritical Trudinger-Moser inequality on the Heisenberg group. The essence of these techniques is to establish the local inequalities on finite domains and then pass them to global inequalities using an argument of level set of functions under consideration. This kind of argument avoids using the Póly-Szegö inequality and thus is also effective in proving the concentration-compactness principle in settings where such symmtrization principle is absent ([10, 11]). This argument has been extensively used by a number of authors and we only refer the reader to see also e.g. [14], [2], [12, 13] for adapations of this circle of ideas.
This sort of ideas has also been used in our proofs of the critical Trudinger-Moser inequality (Theorem 1.3) on complete and noncompact Riemannian manifolds [8]. In the proof of Theorem 1.3 in [8], some clarifications need to be made about the uniformity of the upper bound of the Trudinger-Moser inequality in terms of the bounded domains of the level sets involved. To avoid the complexity, we will then simply give another simple proof by only establishing the uniformity of the bounds of the Trudinger-Moser inequality on geodesic balls within the manifolds (still going through the same proof of Theorem 3.1 in [8] on the geodesic balls or simply applying Theorem 3.1 in [8] on geodesic balls where the Green functions have uniform estimates) together with a Gromov’s covering lemma (similar to the proof of subcritical case in [8]). The proof of Theorem 3.1 in [8] remains valid for geodesic balls with uniform constants for the Trudinger-Moser inequalities for such balls independent of the location of the balls. Since the main purpose of this short note is to clarify the proof of Theorem 1.3 in [8] and in the meantime to give an alternative proof from the level set argument passing from the local inequalities to global one, we have chosen not to give a comprehensive account of the subject (see e.g. [8] for relevant references in the literature).
2. Some useful known lemmas
We recall the critical Moser-Trudinger inequality in [8].
Theorem 2.1.
Let $(M,g)$ be a complete noncompact Riemannian manifold whose Ricci curvature is lower bounded, i.e. $\textit{Ric}\geq\lambda g$ for some constant $\lambda\in\mathbb{R}$. Moreover, we assume that its injectivity radius has a lower bound, i.e. $\textit{inj}(M,g)\geq i>0$ for some constant $i>0$. Then for every $\tau>0$, there exists a constant $C=C(n,\tau,M)$ such that
[…]
Moreover, $\alpha_{n}$ is sharp.
The Ricci lower bound condition guarantees a finite overlapping covering of $M$. This is the so-called Gromov’s covering lemma, which states as follows:
Lemma 2.1 (Hebey [4], Lemma 1.6).
Let $(M,g)$ be a complete Riemannian manifold with Ricci lower bound $\lambda$, let $\rho>0$ be given, then there exists a sequence of points $\{x_{j}\}$ on $M$ such that for any $r\geq\rho$:
  1. (1)
    the family $\{B_{r}(x_{j})\}$ is a uniformly locally finite covering of $M$, there exists an upper bound for the overlapping number in terms of $n,\rho,r,\lambda$;
  2. (2)
    for any $i\neq j$, $B_{r/2}(x_{i})\cap B_{r/2}(x_{j})=\emptyset$.
On the other hand, our geometric assumptions guarantee the existence of harmonic coordinates whose harmonic radius is uniformly bounded from below. To be precise, we have
Lemma 2.2 (Hebey [4], Theorem 1.3).
Let $\alpha\in(0,1)$, $Q>1$, $\delta>0$ and $(M,g)$ satisfy our assumptions. $\Omega(\delta)=\{x\in M:\textit{dist}(x,\Omega)<\delta\}$. Then there exists a positive constant $C=C(n,Q,\alpha,\delta,i,\lambda)$ such that for any $x\in\Omega$, the $C^{0,\alpha}$ harmonic radius $r_{H}(Q,0,\alpha)(x)\geq C$.
Combining Lemma 2.1 and Lemma 2.2, we obtain a sequence of balls $\{B_{\delta}(x_{j})\}$ satisfies following:
  • For $Q>1$ fixed (close to 1), $\delta\leq r_{H}(Q,0,\alpha)$ and $Q^{-1}\delta_{ij}\leq g_{ij}\leq Q\delta_{ij}$ and moreover, $g_{ij}$ satisfies the $C^{0,\alpha}$ estimate:
    $r_{H}^{\alpha}\sup_{y\neq z}\frac{|g_{ij}(z)-g_{ij}(y)|}{d(y,z)^{\alpha}}\leq Q-1.$
  • $\{B_{\delta/2}(x_{j})\}$ covers $M$ and when $i\neq j$, $B_{\delta/4}(x_{i})\cap B_{\delta/4}(x_{j})=\emptyset$;
  • There exists $N=N(n,\lambda,\delta)$ such that any $x\in M$ belongs to at most $N$ of $B_{\delta}(x_{j})$’s.
3. Proof of Theorem 2.1
To give another proof of Theorem 2.1, we achieve this through 2 steps: establishing the uniform Trudinger-Moser inequality on each $B_{\delta}(x_{j})$; gluing all inequalities from Step 1 so we extend our inequality from the local inequalities to a global one. We first prove
Theorem 3.1.
Let $B_{\delta}(x_{j})$ be any ball mentioned above, then there exists a constant $C=C(n,\lambda,i)$ such that
[…]
The proof of Theorem 3.1 consists of 2 steps:
Step 1: Since $u\in W^{1,n}_{0}(B_{\delta}(x_{j}))$, we recall the pointwise representation formula of $u(x)$ in terms of the Dirichlet Green’s function:
$|u(x)|\leq\int_{B_{\delta}(x_{j})}|\nabla G(x,y)||\nabla u(y)|dV_{g}(y).$
The existence of the Green’s function comes from a classical paramatrix argument (see Aubin, Nonlinear Analysis on Manifolds. Monge-Ampére Equations, 1982, Chapter 4). To be precise, for some positive integer $k>\frac{n}{2}$,
$G(x,y)=H(x,y)+\sum_{l=1}^{k}\int_{B_{\delta}(x_{j})}\Gamma_{l}(x,z)H(z,y)dV_{g}(z)+F(x,y),$
where $H(x,y)=\frac{1}{(n-2)\omega_{n-1}}d(x,y)^{2-n}f(d(x,y))$ is the fundamental solution of Laplacian multiplied by a smooth cutoff function $f$, which is identically equal to 1 near diagonal and vanishes when $d(x,y)\geq\textit{inj}(M,g)$. $\Gamma(x,y)=\Gamma_{1}(x,y)=-\Delta_{y}H(x,y)$ and we inductively define
$\Gamma_{l+1}(x,y)=\int_{B_{\delta}(x_{j})}\Gamma_{l}(x,z)\Gamma(z,y)dV_{g}(z).$
The error term $F$ satisfies the equation
(3.1) $\begin{cases}\Delta_{y}F(x,y)=\Gamma_{k+1}(x,y),\\ \ F(x,y)=0\ \ \ \textit{for }y\in\partial B_{\delta}(x_{j}).\end{cases}$
We claim that
(3.2) $|\nabla_{y}G(x,y)|\leq\frac{1}{\omega_{n-1}}d(x,y)^{1-n}(1+Cd(x,y)),$
where $C=C(n,\lambda,Q,\delta)$. Indeed, from the definition of $\Gamma_{l}(x,y)$, their singularity decrease as $l$ grows hence we first look at
$\int_{B_{\delta}(x_{j})}\Gamma(x,z)H(z,y)dV_{g}(z).$
From the asymptotic behavior near singularity, we need to prove the following lemma:
Lemma 3.1.
For any $\alpha,\beta>0$ such that $\alpha+\beta<n$, there exists a constant $C=C(n,\alpha,\beta,Q)$ such that
$\int_{B_{\delta}(x_{j})}d(x,z)^{\alpha-n}d(z,y)^{\beta-n}dV_{g}(z)\leq Cd(x,y)^{\alpha+\beta-n}.$
Proof.
Let $\rho=\frac{d(x,y)}{2}$. Then
(3.3) $\int_{B_{\delta}(x_{j})}d(x,z)^{\alpha-n}d(z,y)^{\beta-n}dV_{g}(z)$
(3.4) $=$ $\left(\int_{B_{\rho}(x)\cap B_{\delta}(x_{j})}+\int_{(B_{3\rho}(y)\setminus B_{\rho}(x))\cap B_{\delta}(x_{j})}+\int_{B_{\delta}(x_{j})\setminus B_{3\rho}(y)}\right)d(x,z)^{\alpha-n}d(z,y)^{\beta-n}dV_{g}(z)$
For the first term of the right hand side,
$\int_{B_{\rho}(x)\cap B_{\delta}(x_{j})}d(x,z)^{\alpha-n}d(z,y)^{\beta-n}dV_{g}(z)$
$\leq$ $\rho^{\beta-n}\int_{B_{\rho}(x)\cap B_{\delta}(x_{j})}d(x,z)^{\alpha-n}dV_{g}(z),$
$\leq$ $Q^{n/2}\omega_{n-1}\rho^{\beta-n}\int_{0}^{\rho}r^{\alpha-n}r^{n-1}dr,$
$\leq$ $Q^{n/2}\omega_{n-1}\rho^{\alpha+\beta-n},$
where in the third line we use the property of harmonic coordinates $Q^{-1}\delta_{ij}\leq g_{ij}\leq Q\delta_{ij}$. Similarly we can estimate the second term of the right hand side of (3.3). For the third term, we have
(3.5) $\int_{B_{\delta}(x_{j})\setminus B_{3\rho}(y)}d(x,z)^{\alpha-n}d(z,y)^{\beta-n}dV_{g}(z)$
(3.6) $\leq$ $C_{n}\int_{B_{\delta}(x_{j})\setminus B_{3\rho}(y)}d(z,y)^{\alpha+\beta-2n}dV_{g}(z),$
(3.7) $\leq$ $C_{n}Q^{n/2}\omega_{n-1}\int_{3\rho}^{\delta}r^{\alpha+\beta-2n}r^{n-1}dr,$
(3.8) $\leq$ $C_{n,\alpha,\beta}Q^{n/2}\omega_{n-1}((3\rho)^{\alpha+\beta-n}-\delta^{\alpha+\beta-n}),$
(3.9) $\leq$ $C_{n,\alpha,\beta}Q^{n/2}\omega_{n-1}\rho^{\alpha+\beta-n}.$
From Lemma 3.1, $\Gamma_{k+1}$ together with its derivatives, are uniformly bounded by a constant $C=C(n,\lambda,i,Q)$. Therefore, to obtain the uniform estimate for Green’s function, we need to prove the uniform $C^{1}$ estimate of $F(x,y)$.
Lemma 3.2.
Let $F(x,y)$ be a $W^{1,2}_{0}(B_{\delta}(x_{j}))$ solution of the Dirichlet problem (3.1), there exists a positive constant $C=C(n,\lambda,i,Q)$ such that $\nabla_{y}F(x,y)||_{\infty}\leq C$.
Proof.
Since $\Gamma_{k+1}$ in (3.1) has no singularity, it is smooth and hence $F$ is the classical smooth solution. We first establish the $C^{0}$ estimate of $F(x,y)$, this follows from a classical barrier argument: consider the following Dirichlet problem on $B_{\delta}(x_{j})$:
(3.10) $\begin{cases}-\Delta u=f(x),\ \ \ &x\in B_{\delta}(x_{j});\\ u=0,\ \ \ &x\in\partial B_{\delta}(x_{j}).\end{cases}$
We are to show: there exists a constant $C=C(n,Q,\delta)$ such that
[…]
In fact, let w(x)=u(x)z(x)w(x)=u(x)-z(x), where
[…]
$x$ is the position variable in terms of harmonic coordinates, and the constant $A$ is left to be chosen. It is easy to verify that
[…]
Since with harmonic coordinates, we have
$\Delta|x|^{2}=2|\nabla x|^{2}+x\Delta x\in[2nQ^{-1},2nQ],$
then we can set $A=2nQ^{-1}$ and hence $\Delta w\leq 0$. Since $w\leq 0$ on $\partial B_{\delta}(x_{j})$, by maximum principle, we have
[…]
Repeat the above argument for u(x)-u(x) then we have
[…]
Therefore we obtain the $C^{0}$ estimate of $F$. Now we recall the Schauder interior estimate of elliptic equations (see Gilbarg-Trudinger [3], Theorem 6.2):
Theorem 3.2 (Schauder estimates).
Let $u\in C^{2,\alpha}(\Omega)$ be a bounded solution in $\Omega$ and the equation
$a^{ij}\partial_{ij}u+b^{i}\partial_{i}u+cu=f,$
where $f\in C^{\alpha}$ and there are positive constants $C_{1},C_{2}$ such that $a^{ij}\xi_{i}\xi_{j}\geq C_{1}|\xi|^{2}$ and
$|a^{ij}|_{C^{0,\alpha}},\ |b^{i}|^{(1)}_{C^{0,\alpha}},\ |c|^{(2)}_{C^{0,\alpha}}\leq C_{2}.$
[…]
In our case, $b^{i},c$ vanish and from Theorem 2.2 in harmonic coordinates $a^{ij}=g^{ij}$ belongs to $C^{0,\alpha}$. $f$ corresponds to $\Gamma_{k}$, which is proved to be uniformly bounded from Lemma 3.1. Therefore we have $|F|^{*}_{C^{2,\alpha}}\leq C(n,\lambda,i,Q,\alpha)$. In specific, $\nabla_{y}F(x,y)||_{\infty}\leq C(n,\lambda,i,Q)$ and we complete the proof. ∎
Step 2: We use Adams’ method to obtain the equivalent version of Theorem 3.1:
Theorem 3.3.
Let $B_{\delta}(x_{j})$ be defined as previous and $T$ be the convolution operator such that
$Tu(x)=\int_{B_{\delta}(x_{j})}|\nabla G(x,y)||\nabla u(y)|dV_{g}(y).$
Then there exists a constant $C=C(n,\lambda,i,Q,\textit{Vol}(B_{\delta}(x_{j})))$, which is increasing with respect to $\textit{Vol}(B_{\delta}(x_{j}))$, such that
[…]
The proof of Theorem 3.3 is exactly the same as that of Theorem 3.1 in [8] which is one of the main results. The parameters $\lambda,i$ do not explicitly appear in the estimates, but will determine $Q$ and $r_{H}(\alpha,0,Q)$, which gives the bounds of $\delta$. This shows the constants in the Moser-Trudinger inequality on any geodesic balls only depend on the radius $\delta$ which in turn depends on $\lambda$ and $i$.
Proof of Theorem 2.1.
For each $j$, we take cutoff function $\psi_{j}\in C_{0}^{\infty}(B_{\delta}(x_{j}))$ satisfying $0\leq\psi_{j}\leq 1$, $\psi_{j}=1$ on $B_{\frac{\delta}{2}}(x_{j})$ and $|\nabla\psi_{j}|\leq\frac{4}{\delta}$. Then $|\nabla\psi_{j}^{2}|=2\psi_{j}|\nabla\psi_{j}|\leq\frac{8}{\delta}\psi_{j}$ and $1\leq\sum_{j}\psi_{j}(x)\leq N$. Thus we have
(3.11) $\int_{M}\phi_{n}(\alpha_{n}|u|^{\frac{n}{n-1}})dV_{g}$
(3.12) $\leq$ $\sum_{j}\int_{B_{\delta/2}(x_{j})}\phi_{n}(\alpha_{n}|u|^{\frac{n}{n-1}})dV_{g},$
(3.13) $\leq$ $\sum_{j}\int_{B_{\delta}(x_{j})}\phi_{n}(\alpha_{n}|\psi_{j}^{2}u|^{\frac{n}{n-1}})dV_{g},$
[…]
(3.16) $\leq$ $C(n,\lambda,i)2^{n}\sum_{j}\int_{M}\psi_{j}^{2n}|\nabla u|^{n}+|\nabla\psi_{j}^{2}|^{n}|u|^{n}dV_{g},$
(3.17) $\leq$ $C(n,\lambda,i)2^{n}\left(N\int_{M}|\nabla u|^{n}dV_{g}+\frac{8^{n}}{\delta^{n}}N\int_{M}|u|^{n}dV_{g}\right),$
(3.18) $\leq$ $C(n,\lambda,i)(N+N\frac{8^{n}}{\delta^{n}\tau^{n}}).$
This is a constant only depending on the initial geometry of the manifold.