### v2

# Optimal Trudinger-Moser inequalities on complete noncompact Riemannian manifolds: Revisit of the argument from the local inequalities to global ones

Abstract.
The main purpose of this short note, on the one hand, is to clarify some part of the proof of Theorem 1.3 in [10] using the level set argument in a simple way, and on the other hand, to give an alternative argument from local inequalities to global ones by using Gromov’s covering lemma. Moreover, we show that the Green functions for the Dirichlet problem on any small geodesic balls (or harmonic coordinate charts) on Riemannian manifolds with positive injective radius and Ricci lower bound have uniform estimates independent of the locations of the balls (or the charts). This will allow us to conclude that the Trudinger-Moser inequality on any level set is uniformly bounded independent of the location of the level set and thus rigorize the proof of Theorem 1.3 in [10]. This latter result of Green’s function estimates is of its independent interest and can be found useful elsewhere.
1. Introduction
The classical Pólya-Szegö symmetrization inequality plays a fundamental role in establishing geometric and functional inequalities in the Euclidean space. However, this symmetrization principle in general fails in many other non-Euclidean settings such as the Heisenberg group or Riemannian manifolds as well as on higher order Sobolev spaces even in the Euclidean space.
To circumvent this obstacle, Lam and Lu [7, 8] developed a symmetrization-free method to prove critical global Trudinger-Moser inequalities on the entire Heisenberg group, and Adams inequalities on the higher order Sobolev spaces on the entire Euclidean spaces. Lam, Lu and Tang [9] establish the subcritical Trudinger-Moser inequality on the Heisenberg group. The essence of these techniques is to establish the local inequalities on finite domains and then pass them to global inequalities using an argument of level set of functions under consideration. This kind of argument avoids using the Póly-Szegö inequality and thus is also effective in proving the concentration-compactness principle in settings where such symmtrization principle is absent ([12, 13]). This argument has been extensively used by a number of authors and we only refer the reader to see also e.g. [16], [4], [14, 15] for adapations of this circle of ideas.
This sort of ideas has also been used in our proofs of the critical Trudinger-Moser inequality (Theorem 1.3) on complete and noncompact Riemannian manifolds [10]. To be precise, we established
Theorem 1.1.
Let $(M,g)$ be a complete noncompact Riemannian manifold whose Ricci curvature is lower bounded, i.e. $\textit{Ric}\geq\lambda g$ for some constant $\lambda\in\mathbb{R}$. Moreover, we assume that its injectivity radius has a lower bound, i.e. $\textit{inj}(M,g)\geq i>0$ for some constant $i>0$. Then for every $\tau>0$, there exists a constant $C=C(n,\tau,M)$ such that
[…]
Moreover, $\alpha_{n}$ is sharp.
In the proof of Theorem 1.1 some clarifications need to be made about the uniformity of the upper bound of the Trudinger-Moser inequality in terms of the bounded domains of the level sets involved. In this note, we will then simply give another simple proof by only establishing the uniformity of the bounds of the Trudinger-Moser inequality on geodesic balls within the manifolds (still going through the same proof of Theorem 3.1 in [10] on the geodesic balls or simply applying Theorem 3.1 in [10] on geodesic balls where the Green functions have uniform estimates) together with a Gromov’s covering lemma (similar to the proof of subcritical case in [10]). The proof of Theorem 3.1 in [10] remains valid for geodesic balls with uniform constants for the Trudinger-Moser inequalities for such balls independent of the location of the balls.
The main result of this note is that the Green functions for the Dirichlet problem on any small geodesic balls on Riemannian manifolds with positive injective radius and Ricci lower bound have uniform estimates independent of the locations of the balls. This will allow us to conclude that the Trudinger-Moser inequality on any level set is uniformly bounded independent of the location of the level set. More precisely, in the proof of Theorem 1.3 in [10], the level set $\Omega(u)$ can be covered by a finite number of small geodesic balls (or harmonic coordinate charts) and the number of such balls (or charts) is uniformly bounded. Therefore, the same level argument from local inequalities to the global ones in [10] are justified.
Since the main purpose of this short note is to rigorize the proof of Theorem 1.3 in [10] and in the meantime to give an alternative proof from the level set argument passing from the local inequalities to global one, we have chosen not to give a comprehensive account of the subject (see e.g. [10] for relevant references in the literature).
2. Some preliminaries in differential geometry
We first recall the Gromov’s covering lemma:
Lemma 2.1 (Lemma 1.6 of Hebey [6]).
Let $(M,g)$ be a complete Riemannian manifold with Ricci lower bound $\lambda$, let $\delta>0$ be given, then there exists a sequence of points $\{x_{j}\}$ on $M$ such that for any $r\geq\delta$:
  1. (1)
    the family $\{B_{r}(x_{j})\}$ is a uniformly locally finite covering of $M$, there exists an upper bound for the overlapping number in terms of $n,\delta,r,\lambda$;
  2. (2)
    for any $i\neq j$, $B_{r/2}(x_{i})\cap B_{r/2}(x_{j})=\emptyset$.
On the other hand, lower bounds for both Ricci curvature and injectivity radius guarantee the existence of $C^{0,\alpha}$ harmonic coordinates. This is due to Anderson-Cheeger [1] (see also Hebey [6]) and to be precise:
Lemma 2.2.
Let $\alpha\in(0,1)$, $Q>1$, $\delta>0$ and $(M,g)$ satisfy $\textit{Ric}\geq\lambda g$ and $\textit{inj}(M,g)\geq i>0$. Then there exists a positive constant $C=C(n,Q,\alpha,\delta,i,\lambda)$ such that for any $x\in\Omega$, the $C^{0,\alpha}$ harmonic radius $r_{H}(Q,0,\alpha)(x)\geq C$.
Combining Lemma 2.1 and Lemma 2.2, we obtain a sequence of balls $\{B_{\delta}(x_{j})\}$ satisfies following:
  • For $Q>1$ fixed (close to 1), $\delta\leq r_{H}(Q,0,\alpha)$ and $Q^{-1}\delta_{ij}\leq g_{ij}\leq Q\delta_{ij}$ and moreover, $g_{ij}$ satisfies the $C^{0,\alpha}$ estimate:
    $r_{H}^{\alpha}\sup_{y\neq z}\frac{|g_{ij}(z)-g_{ij}(y)|}{d(y,z)^{\alpha}}\leq Q-1.$
  • In the overlapping charts $B_{\delta}(x_{i})\cap B_{\delta}(x_{j})$, the transition function satifies
    $y^{k}\circ(x^{l})^{-1}||_{C^{1,\alpha}}\leq C(n,Q,\lambda,\alpha,\delta,i).$
  • $\{B_{\delta/2}(x_{j})\}$ covers $M$ and when $i\neq j$, $B_{\delta/4}(x_{i})\cap B_{\delta/4}(x_{j})=\emptyset$;
  • There exists $N=N(n,\lambda,\delta)$ such that any $x\in M$ belongs to at most $N$ of $B_{\delta}(x_{j})$’s.
Remark 2.1.
Lemma 2.2 is due to Anderson-Cheeger [1]. The $C^{0,\alpha}$ result was achieved by the existence of $W^{1,p}$- harmonic coordinates and a classical Sobolev embedding for $\alpha=1-\frac{n}{p}$. This means in each local chart, the metric tensor satisfies
[…]
Similar embedding also appears for the coordinate functions $x^{i},i=1,\dots,n$ and we have $W^{2,p}$ uniform bounds.
3. Green’s function estimate
For $u\in W^{1,n}_{0}(B_{\delta}(x_{j}))$, we recall the pointwise representation formula of $u(x)$ in terms of the Dirichlet Green’s function:
$|u(x)|\leq\int_{B_{\delta}(x_{j})}|\nabla G(x,y)||\nabla u(y)|dV_{g}(y).$
The existence of the Green’s function comes from a classical paramatrix argument (see e.g. Chapter 4 of Aubin [2]). To be precise, for some positive integer $k>\frac{n}{2}$,
$G(x,y)=H(x,y)+\sum_{l=1}^{k}\int_{B_{\delta}(x_{j})}\Gamma_{l}(x,z)H(z,y)dV_{g}(z)+F(x,y),$
where $H(x,y)=\frac{1}{(n-2)\omega_{n-1}}d(x,y)^{2-n}f(d(x,y))$ is the fundamental solution of Laplacian multiplied by a smooth cutoff function $f$, which is identically equal to 1 near diagonal and vanishes when $d(x,y)\geq\textit{inj}(M,g)$. $\Gamma(x,y)=\Gamma_{1}(x,y)=-\Delta_{y}H(x,y)$ and we inductively define
$\Gamma_{l+1}(x,y)=\int_{B_{\delta}(x_{j})}\Gamma_{l}(x,z)\Gamma(z,y)dV_{g}(z).$
The error term $F$ satisfies the equation
(3.1) $\begin{cases}\Delta_{y}F(x,y)=\Gamma_{k+1}(x,y),\\ \ F(x,y)=0\ \ \ \textit{for }y\in\partial B_{\delta}(x_{j}).\end{cases}$
We claim that
(3.2) $|\nabla_{y}G(x,y)|\leq\frac{1}{\omega_{n-1}}d(x,y)^{1-n}(1+Cd(x,y)),$
where $C=C(n,\lambda,Q,\lambda,\delta)$. Indeed, from the definition of $\Gamma_{l}(x,y)$, their singularity decrease as $l$ grows hence we first look at
$\int_{B_{\delta}(x_{j})}\Gamma(x,z)H(z,y)dV_{g}(z).$
From the asymptotic behavior near singularity, we need to prove the following lemma:
Lemma 3.1.
For any $\alpha,\beta>0$ such that $\alpha+\beta<n$, there exists a constant $C=C(n,\alpha,\beta,Q)$ such that
$\int_{B_{\delta}(x_{j})}d(x,z)^{\alpha-n}d(z,y)^{\beta-n}dV_{g}(z)\leq Cd(x,y)^{\alpha+\beta-n}.$
Proof.
Let $\rho=\frac{d(x,y)}{2}$. Then
(3.3) $\int_{B_{\delta}(x_{j})}d(x,z)^{\alpha-n}d(z,y)^{\beta-n}dV_{g}(z)$
(3.4) $=$ $\left(\int_{B_{\rho}(x)\cap B_{\delta}(x_{j})}+\int_{(B_{3\rho}(y)\setminus B_{\rho}(x))\cap B_{\delta}(x_{j})}+\int_{B_{\delta}(x_{j})\setminus B_{3\rho}(y)}\right)d(x,z)^{\alpha-n}d(z,y)^{\beta-n}dV_{g}(z)$
It is well known that in polar coordinates
$dV=r^{n-1}\left(1-\frac{\textit{Ric}(\theta)}{3}r^{2}+o(r^{2})\right)d\theta dr.$
When $\delta$ is small, providing Ricci curvature lower bounded, the first term of the right hand side,
$\int_{B_{\rho}(x)\cap B_{\delta}(x_{j})}d(x,z)^{\alpha-n}d(z,y)^{\beta-n}dV_{g}(z)$
$\leq$ $\rho^{\beta-n}\int_{B_{\rho}(x)\cap B_{\delta}(x_{j})}d(x,z)^{\alpha-n}dV_{g}(z),$
$\leq$ $C(\lambda,\delta)\omega_{n-1}\rho^{\beta-n}\int_{0}^{\rho}r^{\alpha-n}r^{n-1}dr,$
$\leq$ $C(\lambda,\delta)\omega_{n-1}\rho^{\alpha+\beta-n},$
Similarly we can estimate the second term of the right hand side of (3.3). For the third term, we have
(3.5) $\int_{B_{\delta}(x_{j})\setminus B_{3\rho}(y)}d(x,z)^{\alpha-n}d(z,y)^{\beta-n}dV_{g}(z)$
(3.6) $\leq$ $C_{n}\int_{B_{\delta}(x_{j})\setminus B_{3\rho}(y)}d(z,y)^{\alpha+\beta-2n}dV_{g}(z),$
(3.7) $\leq$ $C_{n}C(\lambda,\delta)\omega_{n-1}\int_{3\rho}^{\delta}r^{\alpha+\beta-2n}r^{n-1}dr,$
(3.8) $\leq$ $C_{n,\alpha,\beta,\lambda,\delta}\omega_{n-1}((3\rho)^{\alpha+\beta-n}-\delta^{\alpha+\beta-n}),$
(3.9) $\leq$ $C_{n,\alpha,\beta,\lambda,\delta}\omega_{n-1}\rho^{\alpha+\beta-n}.$
It is easy to verify that in polar coordinates,
[…]
where $r=d(x,y)$ and $|\widetilde{g}|=\textit{det}(\widetilde{g}_{ij})$ denotes the determinant of the corresponding metric tensor. By carefully choosing the smooth cut-off function $f$, both the first two terms $|(n-3)f^{\prime}-rf^{\prime\prime}|$ and the coefficient of the third term $((n-2)f-rf^{\prime})$ can be uniformly bounded by $C(n,\delta,i)$. With the help of Lemma 3.1, we only need to consider the integral of the form
$\int_{B_{\delta}(x_{j})}|\nabla_{z}\ln\sqrt{|\widetilde{g}|}|(z)H(z,y)dV_{z}.$
An easy change of variablse from polar coordinates to harmonic coordinates shows
$\widetilde{g}_{ij}=g_{kl}\frac{\partial x^{k}}{\partial\widetilde{x}^{i}}\frac{\partial x^{l}}{\partial\widetilde{x}^{j}},$
where $\{x^{k}\}$ and $\{\widetilde{x}^{i}\}$ denotes the coordinate variables respectively. For simplicity we denote the transition matrix $A=\left(\frac{\partial x^{k}}{\partial\widetilde{x}^{i}}\right)$ then
$|\nabla\ln\sqrt{|\widetilde{g}|}|\leq\frac{|\nabla|g||}{2|g|}+\frac{|\nabla|A||}{|A|}.$
Since in harmonic coordinates $Q^{-1}\delta_{ij}\leq g_{ij}\leq Q\delta_{ij}$, we have
$\left|\int_{B_{\delta}(x_{j})}\frac{|\nabla|g||}{|g|}(z)H(z,y)dV_{z}\right|$ $\leq C(n,Q)\int_{B_{\delta}(x_{j})}\frac{|\nabla g(z)|}{d(z,y)^{n-2}}dV_{z}$
Similar to the argument in the proof of Lemma 3.1, the integral can be splitted into 3 pieces. Here we just check the part over $B_{\rho}(x)\cap B_{\delta}(x_{j})$, the other two parts can be treated in a similar manner:
$\int_{B_{\rho}(x)\cap B_{\delta}(x_{j})}\frac{|\nabla g(z)|}{d(z,y)^{n-2}}dV_{z}$ $\leq\frac{C(n)}{d(x,y)^{n-2}}\int_{B_{\rho}(x)\cap B_{\delta}(x_{j})}|\nabla g|dV_{z};$
[…]
$\leq\frac{C(n,Q)}{d(x,y)^{n-2-\frac{n}{p^{\prime}}}}.$
On the other hand, $Q^{-1}\delta_{ij}\leq g_{ij}\leq Q\delta_{ij}$ implies there exists $C(n,Q)>0$ such that $|A|\geq C(n,Q)>0$. Hence we have
$\left|\int_{B_{\delta}(x_{j})}\frac{|\nabla|A||}{|A|}(z)H(z,y)dV_{z}\right|\leq C(n,Q)\int_{B_{\delta}(x_{j})}\frac{|\nabla|A|(z)|}{d(z,y)^{n-2}}dV_{z}$
We can estimate the above integral in a similar manner, but this time we make use of the $W^{2,p}$ estimate of the harmonic coordinate function $x^{i},i=1,\dots,n$. Indeed, in the derivation of the existence of harmonic coordinates, the $L^{p}$ theory of elliptic operators gives the uniform bound of $x^{i}||_{W^{2,p}}$, where $p$ satisfies $\alpha=1-\frac{n}{p}$. Hence there exists $C(n,\lambda,i,Q)$ such that $||\nabla A|_{L^{p}}\leq C(n,\lambda,i,Q)$.
Therefore from Lemma 3.1, $\Gamma_{k+1}$ together with its derivatives, are uniformly bounded by a constant $C=C(n,\lambda,i,Q)$. To obtain the uniform estimate for Green’s function, it suffices to prove the uniform $C^{1}$ estimate of $F(x,y)$.
Lemma 3.2.
Let $F(x,y)$ be a $W^{1,2}_{0}(B_{\delta}(x_{j}))$ solution of the Dirichlet problem (3.1), there exists a positive constant $C=C(n,\lambda,i,Q)$ such that $\nabla_{y}F(x,y)||_{\infty}\leq C$.
Proof.
Since $\Gamma_{k+1}$ in (3.1) has no singularity, it is smooth and hence $F$ is the classical smooth solution. We first establish the $C^{0}$ estimate of $F(x,y)$, this follows from a classical barrier argument: consider the following Dirichlet problem on $B_{\delta}(x_{j})$:
(3.10) $\begin{cases}-\Delta u=f(x),\ \ \ &x\in B_{\delta}(x_{j});\\ u=0,\ \ \ &x\in\partial B_{\delta}(x_{j}).\end{cases}$
We are to show: there exists a constant $C=C(n,Q,\delta)$ such that
[…]
In fact, let w(x)=u(x)z(x)w(x)=u(x)-z(x), where
[…]
$x$ is the position variable in terms of harmonic coordinates, and the constant $A$ is left to be chosen. It is easy to verify that
[…]
Since with harmonic coordinates, we have
$\Delta|x|^{2}=2|\nabla x|^{2}+x\Delta x\in[2nQ^{-1},2nQ],$
then we can set $A=2nQ^{-1}$ and hence $\Delta w\leq 0$. Since $w\leq 0$ on $\partial B_{\delta}(x_{j})$, by maximum principle, we have
[…]
Repeat the above argument for u(x)-u(x) then we have
[…]
Therefore we obtain the $C^{0}$ estimate of $F$. Now we proceed the $C^{1,\alpha}$ Schauder estimate on $B_{\delta}(x_{j})$ so we obtain the $L^{\infty}$ estimate of $\nabla_{y}F(x,y)$. The proof reduces the problem to Euclidean boundary $C^{1,\alpha}$ estimates via harmonic coordinates and a flattening of the boundary. The argument is classical and standard, we only sketch the proof.
Step 1: Reduction to a Euclidean domain. Let $\Phi:B_{r}(p)\to\Omega\subset\mathbb{R}^{n}$ be the harmonic coordinate chart. Let $\widetilde{u}=u\circ\Phi^{-1}$, $\widetilde{f}=f\circ\Phi^{-1}$, and $\widetilde{\varphi}=\varphi\circ\Phi^{-1}$. Then $\widetilde{u}$ solves a uniformly elliptic equation in divergence form on $\Omega$:
[…]
with boundary data $\widetilde{u}=\widetilde{\varphi}$ on $\partial\Omega$. The coefficients satisfy
$\widetilde{a}^{ij},\widetilde{b}^{i},\widetilde{c}\in C^{\alpha}(\overline{\Omega}),\qquad\|\widetilde{a}^{ij}\|_{C^{\alpha}}+\|\widetilde{b}^{i}\|_{C^{\alpha}}+\|\widetilde{c}\|_{C^{\alpha}}\leq C(n,K,i_{0},r).$
Step 2: Flattening the boundary. Since $r<i_{0}$, the boundary $\partial B_{r}(p)$ is smooth, and thus $\partial\Omega$ is $C^{1,\alpha}$. Cover $\partial\Omega$ by finitely many coordinate patches $\{U_{k}\}$ such that each patch is mapped by a $C^{1,\alpha}$-diffeomorphism $\Psi_{k}$ to a half-ball $B_{1}^{+}\subset\mathbb{R}^{n}$. Let $\{\eta_{k}\}$ be a partition of unity subordinate to this cover.
Step 3: Apply Euclidean boundary $C^{1,\alpha}$ estimates. Fix one patch $U_{k}$ and define $v_{k}=\widetilde{u}\circ\Psi_{k}^{-1}$ on $B_{1}^{+}$. Then $v_{k}$ solves a uniformly elliptic equation in divergence form with $C^{\alpha}$ coefficients and boundary data $v_{k}=\varphi_{k}$ on $\partial B_{1}^{+}$. By the Euclidean boundary $C^{1,\alpha}$ estimate (Gilbarg–Trudinger [5]), we have
$\|v_{k}\|_{C^{1,\alpha}(\overline{B_{1/2}^{+}})}\leq C\Big(\|v_{k}\|_{L^{\infty}(B_{1}^{+})}+\|\widetilde{f}\|_{C^{\alpha}(B_{1}^{+})}+\|\varphi_{k}\|_{C^{1,\alpha}(\partial B_{1}^{+})}\Big).$
Step 4: Patch local estimates to the whole domain. Multiplying by $\eta_{k}$ and summing over $k$, we obtain a global estimate on $\Omega$:
$\|\widetilde{u}\|_{C^{1,\alpha}(\overline{\Omega})}\leq C\Big(\|\widetilde{u}\|_{L^{\infty}(\Omega)}+\|\widetilde{f}\|_{C^{\alpha}(\Omega)}+\|\widetilde{\varphi}\|_{C^{1,\alpha}(\partial\Omega)}\Big).$
Using the uniform $C^{1,\alpha}$ control of the coordinate chart $\Phi$, we obtain the desired estimate.
4. Proof of Theorem 1.1
To prove Theorem 1.1, we achieve this through 2 steps: establishing the uniform Trudinger-Moser inequality on each $B_{\delta}(x_{j})$; gluing all inequalities from Step 1 so we extend our inequality from the local inequalities to a global one. We first prove
Theorem 4.1.
Let $B_{\delta}(x_{j})$ be any ball mentioned above, then there exists a constant $C=C(n,\lambda,i)$ such that
[…]
We use Adams’ method to obtain an equivalent version of Theorem 4.1:
Theorem 4.2.
Let $B_{\delta}(x_{j})$ be defined as previous and $T$ be the convolution operator such that
$Tu(x)=\int_{B_{\delta}(x_{j})}|\nabla G(x,y)||\nabla u(y)|dV_{g}(y).$
Then there exists a constant $C=C(n,\lambda,i,Q,\textit{Vol}(B_{\delta}(x_{j})))$, which is increasing with respect to $\textit{Vol}(B_{\delta}(x_{j}))$, such that
[…]
With the help of Green’s function estimates in the last section, the proof of Theorem 4.2 is exactly the same as that of Theorem 3.1 in [10] which is one of the main results. The parameters $\lambda,i$ do not explicitly appear in the estimates, but will determine $Q$ and $r_{H}(\alpha,0,Q)$, which gives the bounds of $\delta$. This shows the constants in the Moser-Trudinger inequality on any geodesic balls only depend on the radius $\delta$ which in turn depends on $\lambda$ and $i$.
Proof of Theorem 1.1.
For each $j$, we take cutoff function $\psi_{j}\in C_{0}^{\infty}(B_{\delta}(x_{j}))$ satisfying $0\leq\psi_{j}\leq 1$, $\psi_{j}=1$ on $B_{\frac{\delta}{2}}(x_{j})$ and $|\nabla\psi_{j}|\leq\frac{4}{\delta}$. Then $|\nabla\psi_{j}^{2}|=2\psi_{j}|\nabla\psi_{j}|\leq\frac{8}{\delta}\psi_{j}$ and $1\leq\sum_{j}\psi_{j}(x)\leq N$. Thus we have
(4.1) $\int_{M}\phi_{n}(\alpha_{n}|u|^{\frac{n}{n-1}})dV_{g}$
(4.2) $\leq$ $\sum_{j}\int_{B_{\delta/2}(x_{j})}\phi_{n}(\alpha_{n}|u|^{\frac{n}{n-1}})dV_{g},$
(4.3) $\leq$ $\sum_{j}\int_{B_{\delta}(x_{j})}\phi_{n}(\alpha_{n}|\psi_{j}^{2}u|^{\frac{n}{n-1}})dV_{g},$
[…]
(4.6) $\leq$ $C(n,\lambda,i)2^{n}\sum_{j}\int_{M}\psi_{j}^{2n}|\nabla u|^{n}+|\nabla\psi_{j}^{2}|^{n}|u|^{n}dV_{g},$
(4.7) $\leq$ $C(n,\lambda,i)2^{n}\left(N\int_{M}|\nabla u|^{n}dV_{g}+\frac{8^{n}}{\delta^{n}}N\int_{M}|u|^{n}dV_{g}\right),$
(4.8) $\leq$ $C(n,\lambda,i)(N+N\frac{8^{n}}{\delta^{n}\tau^{n}}).$
This is a constant only depending on the initial geometry of the manifold.
Here we make some remarks about our alternative proof.
Remark 4.1.
To establish global inequalities from local ones, we choose appropriate cutoff functions $\psi_{j}$ such that $\nabla(\psi_{j}^{2}u)||_{L^{n}}\leq 1$. Thus the scale parameter $\delta$ actually depends on $\tau$. On the other hand, in the Green’s function estimates, we make use of both $W^{1,p}$ bounds of $g$ and $W^{2,p}$ bounds of $x^{i}$. There we need to choose proper value of $p$, hence $\alpha$. This can be done since the harmonic coordinates exists for arbitrary chosen $\delta$’s and $\alpha$’s. In other words, our inequalities can only be proved in specific harmonic coordinates.
Remark 4.2.
There is another version of harmonic coordinates one can use to prove our main theorem. In [1], it reads as a collection of finitely overlapping charts $\{U_{j},F_{j}\}$, where $F_{j}:U_{j}\to\mathbb{R}^{n}$ is of the form $U_{j}=F_{j}^{-1}(B_{r_{H}}(0))$. In such choice of harmonic coordinates, the a priori estimate of $F(x,y)$ will be significantly simplified since the Dirichlet problem is taken on Euclidean balls whose boundary is smooth and can be uniformly finitely covered. On the other hand, the domains $\{U_{j}\}$ share the same Gromov’s covering properties. Noticing that $\textit{Vol}(U_{j})=\int_{U_{j}}dV=\int_{B_{r_{H}}(0)}\sqrt{|g|}dx\sim C(n,Q)r_{H}^{n}$, it is still possible to perform the local to global argument using this collection of charts, instead of geodesic balls.